Find the norm of the linear functional f defined by

𝑓 π‘₯ =integral -1 to 0π‘₯ 𝑑 𝑑𝑑 βˆ’ integral 0 to 1π‘₯ 𝑑 𝑑𝑑

whereπ‘₯ ∈ [βˆ’1 , 1]


1
Expert's answer
2021-11-30T16:41:52-0500

"|f(x)|=""|\int^0_{-1}x(t)dt-\int^1_{0}x(t)dt|\le |\int^0_{-1}x(t)dt|+|\int^1_{0}x(t)dt|"


"|\int^0_{-1}x(t)dt|+|\int^1_{0}x(t)dt|\le {\sup}_{t\isin [-1,0]}|x(t)|+{\sup}_{t\isin [0,1]}|x(t)|\le 2{\sup}_{t\isin [-1,1]}|x(t)|="


"=2||x(t)||_{\infin}"


thus, "||f||\le2"


if we consider the sequence

"x_n^*(t)=\begin{cases}
 1 &\text{if } -1\le t\le -1/n \\
 -nt &\text{if } -1/n< t< 1/n \\
 -1 &\text{if } 1/n\le t\le 1
\end{cases}"


"x_n^*(t)\isin C[-1,1]" for each "n\isin N"

Then we see that

"||f||=sup_{x\in C[-1,1]:||x||=1}|f(x)|\ge |f(x_n^*)|\forall n \isin N"

since "|f(x_n^*)|=|2+1\/n^2-1\/n|" :

"||f||\ge |2+1/n^2-1/n| \implies ||f|| \ge 2"


so, "||f|| =2"